30 C of charge passes in a 5 second period through a resistive material of cross- sectional area of a circle with radius 1 x 10-3 m. How much of the current goes through the cross-section in A?




1
Expert's answer
2021-12-07T20:12:20-0500

The current is


"I=\frac QT."


The current density is


"J=\frac{I}{A}=\frac {Q}{tA}=\frac {Q}{t\pi r^2},\\\space\\
J=1.91·10^6\text{ A/m}^2."